# Understanding and Imitating Human-Robot Motion with Restricted Visual Fields

### v2

@@ IV Imitating Human Behavior
Figure 4: Participants used a mouse and keyboard to navigate the optimal path for the robot to reach the goal in the bottom right corner. Their vision was kept limited to simulate human perception in real-world scenarios. The obstacles, depicted as asteroids, each followed a unique movement pattern that was unknown to the participants—some moved quickly, others slowly, and some remained stationary. We collected trajectory data from the participants to better understand how humans observe their surroundings and make decisions in this driving-like scenario.
We collect human trajectory data to capture realistic navigation strategies, aiming to predict and understand human actions for improved collaboration. By modeling these behaviors, our approach enables the system to anticipate and adapt to human decision-making, enhancing its ability to operate effectively in dynamic and complex environments. Participants gave informed consent under the oversight of an Institutional Review Board. [^1] Protocol 65022. approved by the Administrative Panels for the Protection of Human Subjects at Stanford University, On February 29, 2024.
@@ IV-A1 Environment Design and Game Mechanics
The game environment consisted of a rectangular arena with a fixed starting point for the robot and a fixed goal location. Participants controlled the robot using standard keyboard inputs, enabling them to move in any direction. The robot’s speed was constrained to mimic realistic movement dynamics, adding further complexity to the task. Participants had to carefully balance speed with accuracy to avoid collisions while reaching the goal efficiently.
@@ IV-A2 Limited Field of View
To simulate real-world constraints, participants’ vision was restricted to a cone-shaped field of view (similar to a flashlight beam). This limited their ability to observe the entire environment at once, as only obstacles within the cone were visible. However, the goal remained visible at all times, giving participants a consistent reference point to navigate toward. The field of view had the following characteristics: a fixed $\theta_{\text{obs}}$ in 45 degrees and fixed $r_{\text{obs}}$.
@@ IV-A3 Data Collection
During the experiment, we recorded two critical types of trajectory data: Agent Trajectory Data: This included the position (x, y) and orientation ($\psi$) of the human-controlled robot at each timestep. This data captured the full path taken by participants as they navigated toward the goal. Obstacle Trajectory Data: Both static and moving obstacles had their positions recorded throughout each trial. Static obstacles maintained constant positions, while dynamic obstacles followed specific motion patterns, which were also logged for later analysis. We used data from 8 participants to ensure a diverse set of navigation behaviors.
@@
Figure 5: Trajectories comparison between human data and model inference. All the trajectories have the same environment within this figure.
@@ IV-B Transformer-Based Diffusion for Navigation
The collected human trajectory data serves as the foundation for training our navigation model, aimed at replicating human-like obstacle avoidance and goal-reaching strategies. We employ a transformer-based diffusion model with multi-head cross-attention to dynamically process the agent’s view and make real-time action predictions. The model uses a multi-head cross-attention mechanism, inspired by the Transformer architecture [27], within the class-conditional U-Net [28]. During testing, noisy action inputs and observation embeddings are processed through each U-Net layer. The cross-attention mechanism allows the model to focus on various aspects of the environment, such as obstacle proximity or goal direction, dynamically adapting to the most relevant features at each timestep.The model outputs the action changes, represented as $\Delta x,\Delta y,\Delta\psi$
@@ IV-C1 Original game environment
The Fréchet distances between the actual and predicted trajectories (as shown in Figure 5) were computed using the method described in the numerical simulations section. The average normalized Fréchet distance was found to be 11.3 ± 0.06, indicating a strong alignment between the model’s trajectory predictions and human trajectories in the original game environment. Additionally, the minimum Fréchet distance observed was 4.48, demonstrating the model’s capability to closely replicate human-like navigation under certain conditions.
@@ IV-C2 Scaled-Down Lab Environment
To demonstrate the agent’s ability to generalize beyond the training environment, we generated four scenarios with randomly positioned obstacles (See Fig. 6). In each scenario, the agent begins at the bottom-left corner and aims to reach the goal located at the top-right corner. These examples illustrate that the agent can adapt its trajectory to avoid obstacles while efficiently navigating toward the goal, dynamically controlling its observation space to prioritize obstacle avoidance in a manner that closely mimics human behavior, thereby showcasing its robustness in handling previously unseen environments.
Figure 6: Trajectories generation in unknown environment.
@@ IV-D Safety Analysis of Trajectories Data
Figure 7: This percentage represents how often the robot came dangerously close to obstacles during its movement. Run 1: Human data from data collection. Run 2: Diffusion policy output trajectories with observation model. Run 3: Diffusion policy output trajectories without observation model
@@ V Hardware Experiment
where $p_{t}$ and $q_{t}$ are $x$ and $y$ coordinates of the positions in the 2D plane, $\phi_{t}$ is the heading angle from positive x-axis, $v_{t}$ is the forward velocity, $\delta_{t}$ is the steering angle, $\psi_{t}$ is the orientation of the sensor skirt, $\Delta\psi_{t}$ is the change in sensor skirt orientation, and $\omega_{t}$ is the steering rate of the agent at time $t$. We employ the transformer diffusion model from Section IV for the bicycle dynamics. Note that output from trained diffusion policy is $\Delta x,\Delta y,\Delta\psi$. We use the following inputs for the bicycle from the diffusion model output
$v_{t}=K_{v}\cdot||\delta x||,$ $\phi_{d}=\tan^{-1}(\Delta x_{2},\Delta x_{1})$
@@
where |||||$\cdot$||| | ⋅ | | represents the Euclidean norm, $\Delta x_{i}$ is the $i^{th}$ component of the vector $\Delta x$, $\phi_{d}$ is the desired heading angle for the robot to move towards the next waypoint, and $K_{v}$ and $K_{\omega}$ are hyperparameters tuned for the car. Furthermore, to make the policy real-time applicable, we use a Denoising Diffusion Implicit Models (DDIM) [29] based approach to speed up the denoising process. Instead of 100 DDPM steps, we use 10 DDIM steps followed by 10 DDPM steps to obtain control inputs for the car. This approach makes the policy applicable to the car in real-time. We use the VICON motion capture system to obtain real-time position updates for the car. We use open-source Robotics Operating System [30] to communicate with the car and send commands.
We employ the trained policy from humans on the car. The experiment results can be found in Figure 2. As can be seen, the car successfully navigates the environment with obstacles. Like humans, the car changes its observation direction to look around, detect obstacles, and avoid them while moving toward the goal in real-time.
@@ VI CONCLUSIONS
In this work, we show that for agents who have limited observation characterized by a limited field of view and viewing distance and who may not observe objects within the field of view, it is important to model these observation limitations in addition to the policy in order to correctly predict the motion and behavior of the agent. We demonstrated the utility of approximating these observation parameters for robots, and we showed that by observing humans navigating a cluttered scene, we could imitate their behavior both for motion and independent observation strategies, and we showed that this could be executed on physical robotic hardware.

### v3

@@ IV Imitating Human Behavior
Figure 4: Participants used a mouse and keyboard to navigate to reach the goal in the bottom right corner. Their vision was kept limited to simulate human perception in real-world scenarios. The obstacles followed a unique movement pattern unknown to the participants. Some obstacles moved quickly, others slowly, and some remained stationary. We collected trajectory data from the participants to better understand how humans observe their surroundings and make decisions in this driving-like scenario.
This section focuses on replicating human behavior using data collected directly from human demonstrations. We collect human trajectory data to capture realistic navigation strategies, aiming to predict and understand human actions for improved collaboration. By modeling these behaviors, our approach enables the system to anticipate and adapt to human decision-making, enhancing its ability to operate effectively in dynamic and complex environments.
@@ IV-A1 Environment Design and Game Mechanics
The game environment consisted of a rectangular area with a fixed starting point for the robot and a fixed goal location. Participants controlled the robot using standard keyboard inputs, enabling them to move in any direction. The participants used the mouse to control the robot’s viewing direction $\psi$, which determined how the field of view rotated within the environment. The robot’s speed was constrained to mimic realistic movement dynamics, adding further complexity to the task. Participants had to carefully balance speed with accuracy to avoid collisions while reaching the goal efficiently. We aim to learn both navigation and camera control strategies from this setup using the collected data to replicate how humans coordinate movement and visual attention during the task.
@@ IV-A2 Limited Field of View
To simulate real-world constraints, participants’ vision was restricted to a cone-shaped field of view. This limited their ability to observe the entire environment at once, as only obstacles within the cone were visible. However, the goal remained visible at all times, giving participants a consistent reference point to navigate toward. The field of view had fixed $\theta_{\text{obs}}$ and fixed $r_{\text{obs}}$.
@@ IV-A3 Data Collection
Participants played the game on their computer and gave informed consent under the oversight of an Institutional Review Board. [^1] Protocol 65022. approved by the Administrative Panels for the Protection of Human Subjects at Stanford University, On February 29, 2024. During the experiment, we recorded two critical types of trajectory data: Agent Trajectory Data: This included the position (x, y) and orientation ($\psi$) of the human-controlled robot at each timestep. This data captured the full path taken by participants as they navigated toward the goal. Obstacle Trajectory Data: Both static and moving obstacles had their positions recorded throughout each trial. Static obstacles maintained constant positions, while dynamic obstacles followed specific motion patterns, which were also logged for later analysis. We used data from 8 participants to ensure a diverse set of navigation behaviors.
@@
Figure 5: Qualitative illustration of the human data and model inference. All the trajectories have the same environment within this figure.
@@ IV-B Transformer-Based Diffusion for Navigation
The collected human trajectory data serves as the foundation for training our navigation model, aimed at replicating human-like obstacle avoidance and goal-reaching strategies. We employ a transformer-based diffusion model that uses a multi-head cross-attention mechanism, inspired by the Transformer architecture [29], with a class-conditional U-Net [30]. During testing, noisy action inputs and observation embeddings are processed through each U-Net layer. The cross-attention mechanism allows the model to focus on various aspects of the environment, such as obstacle proximity or goal direction, dynamically adapting to the most relevant features at each timestep. The model outputs the action changes, represented as $\Delta x,\Delta y,\Delta\psi$.
@@ IV-C1 Original game environment
We computed the Fréchet distances between the actual and predicted trajectories (Figure 5). The average normalized Fréchet distance was found to be 11.3 ± 0.06, indicating a strong alignment between the model’s trajectory predictions and human trajectories in the original game environment. Additionally, the minimum Fréchet distance observed was 4.48, demonstrating the model’s capability to closely replicate human-like navigation under certain conditions.
@@ IV-C2 Scaled-Down Lab Environment
To demonstrate the agent’s ability to generalize beyond the training environment, we generated four scenarios with randomly positioned obstacles (See Fig. 6). In each scenario, the agent begins at the bottom-left corner and aims to reach the goal located at the top-right corner. These examples illustrate that the agent can adapt its trajectory to avoid obstacles while efficiently navigating toward the goal, dynamically controlling its observation space to prioritize obstacle avoidance in a manner that closely mimics human behavior. This behavior showcases the robustness of the policy in handling previously unseen environments.
Figure 6: Trajectory generation in an unknown environment.
@@ IV-D Safety Analysis of Trajectories Data
Figure 7: Comparison of obstacle proximity rates across human and model-generated trajectories. Obstacle proximity rate is defined as the fraction of time steps during which the agent was within 20 units of at least one obstacle. The value of $\mu$ shown above each subplot represents the mean obstacle proximity rate for all trajectories in the corresponding run. Run 1: 150 Human demonstration trajectories. Run 2: 20 trajectories generated by the diffusion policy with the observation model. Run 3: 20 trajectories generated by the diffusion policy without the observation model. All obstacles used in these experiments follow the same movement patterns as those from the human data collection environment. Generated trajectories with the observation model better replicate human navigation behavior, with a lower risk of entering danger zones.
@@ V Hardware Experiment
where $p_{t}$ and $q_{t}$ are $x$ and $y$ coordinates of the positions in the 2D plane, $\phi_{t}$ is the heading angle from positive x-axis, $v_{t}$ is the forward velocity, $\delta_{t}$ is the steering angle, $\psi_{t}$ is the orientation of the sensor skirt, $\Delta\psi_{t}$ is the change in sensor skirt orientation, $\omega_{t}$ is the steering rate of the agent at time $t$, and $L$ is the lenght of the car. We employ the transformer diffusion model from Section IV for the bicycle dynamics. Note that in this case, the output from the trained diffusion policy is $\Delta x,\Delta y,\Delta\psi$. However, since the bicycle model requires linear and angular velocities, we convert the diffusion model outputs ($\Delta x,\Delta y,\Delta\psi$) into the corresponding velocity commands via
[…]
@@
where $||\cdot||$ represents the Euclidean norm, $\Delta x_{i}$ is the $i^{th}$ component of the vector $\Delta x$, $\phi_{d}$ is the desired heading angle for the robot to move towards the next waypoint, and $K_{v}$ and $K_{\omega}$ are hyperparameters tuned for the car. Furthermore, to make the policy inference fast enough for real-time implementation, we use a Denoising Diffusion Implicit Model (DDIM) [31] based approach to speed up the denoising process. Instead of 100 DDPM steps, we use 10 DDIM steps followed by 10 DDPM steps to obtain control inputs for the car. This approach makes the policy applicable to the car in real-time. We use the VICON motion capture system to obtain real-time position updates for the car. We use the Robotics Operating System [32] to communicate with the car and send commands.
We employ the trained policy from humans on the car. The experiment results can be found in Figure 2. As can be seen, the car successfully navigates the environment with obstacles. Like humans, the car changes its observation direction to look around, detect obstacles, and avoid them while moving toward its goal in real-time.
@@ VI CONCLUSIONS
In this work, we show that for agents who have limited perception capabilities characterized by a limited field of view, viewing distance, and failure to detect objects within their field of view, it is important to model these observation limitations in addition to their policy in order to correctly predict the motion and behavior of the agent. We demonstrated the utility of approximating these observation parameters for robots, and we showed that by observing humans navigating a cluttered scene and explicitly reasoning about their limited observation space and observation strategies, we could better imitate their behavior. We further showed that the resulting policy could be executed on physical robotic hardware.